You are given the following income-expenditures model for the economy of Vulcan.

C = 200 + .8Yd

T = 50

G = 100

I = 140

a) What is the equilibrium level of income in Vulcan?

b) At the equilibrium level of income, what is the amount of savings?

c) If government spending increases to 150, what is the new level of equilibrium income?
Answer #1

a) At equilibrium, Y = C + G + I

Y= 200 + 0.8(Y-50) + 100 + 140

0.2Y = 400

Y= 2000

b) C= 200 + 0.8(2000-50)= 1760

Savings = Y-C = 2000-1760= 240

c) multiplier = 1/(1-mpc) = 1/(1-0.8)= 5

Increase in G = 150-100= 50

Increase in equilibrium output = 5*50= 250

New level of equilibrium income = 2000+250= 2250